OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

OpenStudy (anonymous):

\[3\sqrt{2x}+x \sqrt{8x}-5\sqrt{18x}\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@esamalaa

OpenStudy (anonymous):

can you simplify \[\sqrt{8x} and \sqrt{18x}\to a \sqrt{2x}\]

OpenStudy (anonymous):

a=??

OpenStudy (anonymous):

and them make square root 2x as prim term